OpenStudy (anonymous):

Please help me! Differentiate:

OpenStudy (anonymous):

\[y=\frac{ \cosh(\cos2x) }{ \arctan \sqrt{x} }\]

zepdrix (zepdrix):

Woah! This one is a doozy XD

zepdrix (zepdrix):

Hmm doesn't look like we have any nice simplification, so jump right in using the quotient rule.

zepdrix (zepdrix):

\[\large y'=\frac{[\cosh (\cos 2x)]'\arctan \sqrt{x} - \cosh (\cos 2x)[\arctan \sqrt x]'}{(\arctan \sqrt x)^2}\]

OpenStudy (anonymous):

okay!

zepdrix (zepdrix):

So the setup is quite messy... but besides that, we don't have too much work to do. We just need to differentiate the 2 primed terms. Maybe we should do those off to the side. The hyperbolic cosine is going to have a few chains to it. Hmm.

zepdrix (zepdrix):

Remember the derivative of cosh? :D

OpenStudy (anonymous):

yes! sinh

zepdrix (zepdrix):

\[\large [\cosh (\cos 2x)]'=[\sinh (\cos 2x)](\cos 2x)'\] Yah sounds good c: that takes care of the outer function.

OpenStudy (anonymous):

okay!

OpenStudy (anonymous):

I just don't know how to do the derivative of \[\arctan \sqrt{x}\]

zepdrix (zepdrix):

So we're going to stuff it into this form, then apply the chain rule as we usually would.\[\huge (\arctan z)'=\frac{1}{1+z^2}z'\]

zepdrix (zepdrix):

\[\huge (\arctan \sqrt x)'=\frac{1}{1+(\sqrt x)^2}(\sqrt x)'\] Understand what we did there?

OpenStudy (anonymous):

yupp! What would the derivative of \[\sqrt{x}\] be?

zepdrix (zepdrix):

oh comeon morgan! really? -_- lol

zepdrix (zepdrix):

you should have that one memorized by now ^^ hehe If you can't remember, change it to a fractional exponent :D then apply the power rule.

OpenStudy (anonymous):

OH! sorry, i had a brain fart! lol

OpenStudy (anonymous):

would the answer be: \[y \prime=\frac{ 2(\arctan \sqrt{x})(\sinh)(\sin2x)-\cosh(\cos2x)(\frac{ 1 }{ 1+x })(1/2x ^{-1/2}}{ (\arctan \sqrt{x})^{2} }\]

zepdrix (zepdrix):

Hmmm we're close. Your sinh term doesn't have anything written inside of it. :O hmmm

OpenStudy (anonymous):

what do you mean?

zepdrix (zepdrix):

the sinh function should have an inside.. you ripped his guts out, where did they go? :(

OpenStudy (anonymous):

Okay, I get it. My brother just helped me! Lol. Can you help me with the derivative of \[3^{arcsinx}\] ??

zepdrix (zepdrix):

Yay, brother to the rescue! :D

zepdrix (zepdrix):

Remember how to take the derivative of an exponential with base that IS NOT e? Here is a quick refresher just in case. \[\huge (3^x)'=3^x(\ln 3)\]

OpenStudy (anonymous):

i get it!

zepdrix (zepdrix):

Just don't forget to apply the chain rule to the exponent!\[\huge (3^{\arcsin x})'=(3^{\arcsin x})(\ln 3)(\arcsin x)'\]

zepdrix (zepdrix):

Woops that was too big lemme try that again. \[\large (3^{\arcsin x})'=(3^{\arcsin x})(\ln 3)(\arcsin x)'\]